# Thompson Sampling via Fine-Tuning of LLMs

### v1

@@ Abstract
Bayesian optimization in large unstructured discrete spaces is often hindered by the computational cost of maximizing acquisition functions due to the absence of gradients. We propose a scalable alternative based on Thompson sampling that eliminates the need for acquisition function maximization by directly parameterizing the probability that a candidate yields the maximum reward. Our approach, Thompson Sampling via Fine-Tuning (ToSFiT) leverages the prior knowledge embedded in prompt-conditioned large language models, and incrementally adapts them toward the posterior. Theoretically, we derive a novel regret bound for a variational formulation of Thompson Sampling that matches the strong guarantees of its standard counterpart. Our analysis reveals the critical role of careful adaptation to the posterior probability of maximality—a principle that underpins our ToSFiT algorithm. Empirically, we validate our method on three diverse tasks: FAQ response refinement, thermally stable protein search, and quantum circuit design. We demonstrate that online fine-tuning significantly improves sample efficiency, with negligible impact on computational efficiency.
@@ 1 Introduction
Bayesian optimization (Kushner, 1964; Garnett, 2023) is one such algorithmic framework that leverages Bayesian inference for large-scale experimental design and automated discovery, particularly in settings where evaluations are costly or time-consuming. As a strategy for optimizing expensive black-box reward functions, it maintains a posterior distribution over the unknown rewards—typically modeled as a Gaussian process over a domain XX—and uses this model to guide the search for promising configurations. New candidates are selected by maximizing an acquisition function that balances two objectives: exploring uncertain regions to gather new information, and exploiting areas that are already known to perform well (Kushner, 1964; Auer, 2002).
Figure 1: ToSFiT treats candidate generation as Thompson sampling and fine-tunes the LLM toward the posterior PoM to ensure consistency.
Among existing acquisition strategies, Thompson sampling (Thompson, 1933; Russo et al., 2018) stands out due to its state-of-the-art convergence guarantees (Russo & Van Roy, 2014; 2016; Chowdhury & Gopalan, 2017) and its strong empirical performance (Chapelle & Li, 2011), despite being one of the earliest strategies for Bayesian optimization. Thompson sampling draws a function from the reward posterior, which effectively serves as an acquisition function, and selects the point that maximizes it. The resulting evaluation points are distributed according to the probability of maximality (PoM) of rewards (Menet et al., 2025).
@@
In this work, we scale Bayesian optimization to large unstructured discrete spaces by instantiating Thompson Sampling as fine-tuning of large language models (ToSFiT, see Figure 1). Our approach builds on Variational Bayesian Optimistic Sampling (VBOS), established by O’Donoghue & Lattimore (2021), but introduces a key distinction: instead of optimizing toward the posterior probability of maximality starting from a uniform policy, we initialize it with a prompt-conditioned pre-trained LLM, and gradually adapt it toward the posterior. By treating generated proposals as Thompson samples, ToSFiT avoids intractable acquisition function maximization. Our contributions are:
@@
    We validate our method on FAQ response refinement, protein optimization, and quantum circuit design. Across all settings, ToSFiT finds much better solutions than Unguided Generation, Post-Generation TS, Actor Critic and Soft Actor Critic.
@@ 2 Preliminaries
Consider an unknown Gaussian reward vector $R$ over a discrete domain $X$ with mean $\mu$ and covariance $K$, i.e., $R\sim\mathcal{N}(\mu,K)$. Thompson sampling suggests conducting Bayesian optimization by repeatedly conditioning the Gaussian reward vector $R$ on an observation $Y_{x}:=R_{x}+\mathcal{N}(0,\sigma_{n}^{2})$ at a point $x\in X$ sampled from the posterior probability of maximality
@@
In practice, sampling from the probability of maximality is typically implemented by drawing a realization from the reward posterior and then selecting the point that maximizes this sample—effectively treating the realization as an acquisition function (Russo et al. 2018, see Figure 2). However, in large unstructured discrete domains, such maximization becomes intractable. To address this, we propose to directly parameterize the probability of maximality using a generative LLM. This approach, illustrated in Figure 1, avoids explicit maximization and, following the framework of O’Donoghue & Lattimore (2021), ensures consistency with the underlying reward model and thus vanishing regret (see Theorem 1).
@@ 2.2 Regret Bounds for Bayesian Optimization
Beyond the Bayesian setting, regret guarantees can also be established for the model-agnostic setting that assumes the reward $r$ to lie in a reproducing kernel Hilbert space $\mathbb{H}$ induced by a kernel $K:X\times X\to\mathbb{R}_{+}$. [^1] The expressivity of a reproducing kernel Hilbert space can vary widely with the choice of its kernel. For instance, the linear kernel induces the reproducing kernel Hilbert space of (all) linear functions, whereas the radial basis function kernel corresponds to a smooth subset of the square-integrable functions $L_{2}$ that can uniformly approximate any continuous function on a compact set. In this case, a Gaussian reward model is still used for inference, but its amplitude (i.e., prior uncertainty) must be chosen large enough to accommodate the complexity of $r$, as measured by the norm $\|r\|_{K}$. Notable examples of such model-agnostic algorithms include GP-UCB (Srinivas et al., 2010) and GP-TS (Chowdhury & Gopalan, 2017), a variant of Thompson sampling.
@@ 2.3 Variational Bayesian Optimistic Sampling
According to O’Donoghue & Lattimore (2021) and Tarbouriech et al. (2024), Gaussian PoM is effectively approximated by the VBOS policy $\tilde{\pi}$, the distribution that maximizes the functional
@@
Here, the reward surrogate $\hat{r}_{x}:=\mu_{x}+\sqrt{-2\ln({\pi_{x}})}\cdot\sigma_{x}$ represents a high-probability upper bound on $R_{x}$ (O’Donoghue & Lattimore, 2021). VBOS reveals an intimate connection (Tarbouriech et al., 2024; Menet et al., 2025) between PoM and upper confidence bounds augmented with entropy regularization (Ziebart, 2010; Haarnoja et al., 2018). In addition to its variational form, Menet et al. (2025) pointed out a near-closed form with almost-linear runtime in $|X|$, given by
@@ 3 Thompson Sampling via Fine-Tuning
Algorithm 1 ToSFiT
@@ 3.1 Scaling Gaussian Processes
The Moore–Aronszajn theorem states that any kernel $K$ can be expressed as an inner product in a reproducing kernel Hilbert space $\mathbb{H}$, i.e., $K(x,z)=\langle\phi(x),\phi(z)\rangle_{\mathbb{H}}$ for a feature map $\phi:X\to\mathbb{H}$. Thus we assume, without loss of generality, a linear kernel in $\mathbb{H}$, enabling scalable Gaussian process inference. Flexibility is retained through the choice of feature map $\phi$, which can be obtained from fixed embedding models (Ranković & Schwaller, 2023; Wei et al., 2022), learned adaptively during optimization (Ranković & Schwaller, 2025), or designed for the task at hand (Greenhill et al., 2020).
Furthermore, as detailed in Appendix B, we leverage a formulation of linear Gaussian processes that enables conditioning on observations, computing the reward posterior, and performing marginal likelihood maximization—all in closed form. As a result, both the computational and memory complexity scale in $\Theta(\text{dim}(\mathbb{H})^{2})$, independently of the number of observations, and has negligible overhead (see Table 1 in Section C.1 of the Appendix).
3.2 Gradients of VBOS
We derive explicit gradients of the VBOS objective, offering new insights into its optimization landscape. Note that the score trick (Williams, 1992) cannot be applied blindly, since the reward surrogate $\hat{r}_{x}:=\mu_{x}+\sqrt{-2\ln({\pi_{x}^{\theta}})}\cdot\sigma_{x}$ itself depends on the policy $\pi^{\theta}$.
@@ 3.3 Stabilizing Gradients of VBOS
Here, $\hat{\hat{r}}^{\theta}_{x_{i}}-\xi_{i}$ is referred to as the advantage function. In practice, this estimator can suffer from high variance, depending on the choice of baselines $\xi_{i}\in\mathbb{R}$. To address this, we adopt the Reinforce Leave-One-Out (RLOO) baseline (Kool et al., 2019), where each $\xi_{i}$ is set to the average of the other surrogates in the batch: $\xi_{i}=\tfrac{1}{B-1}\sum_{j\not=i}\hat{\hat{r}}^{\theta}_{x_{j}}$. This technique has been demonstrated to outperform more complex alternatives like Proximal Policy Optimization (Schulman et al., 2017) in fine-tuning large language models (Ahmadian et al., 2024). To further stabilize learning, we normalize the advantage function by its empirical standard deviation, derived from the empirical second moments: $\widehat{\text{advantage std}}=\sqrt{\tfrac{1}{B}\sum_{h}(\hat{\hat{r}}^{\theta}_{x_{h}}-\xi_{h})^{2}}$, which amounts to a variance-adaptive learning rate. As shown in Proposition 3 of Appendix A, standardized RLOO is mathematically equivalent to the advantage function used in Group Relative Policy Optimization (Shao et al., 2024).
@@ 4 Theory
Even the exact VBOS policy $\tilde{\pi}$ only conducts approximate Thompson sampling. Nevertheless, O’Donoghue & Lattimore (2021) prove an upper bound on the cumulative regret incurred when sampling a bandit according to exact VBOS, i.e., for $x^{t}\!\sim\!\tilde{\pi}^{t}$. The bound reads $\mathbb{E}[\sum_{t}R^{*}-R_{x^{t}}]\leq\sqrt{2|X|T\ln|X|(1\!+\!\ln T)}$. Note that the structure of the kernel is not taken into account, i.e., the worst-case bandit with independent arms is assumed. In Theorem 1, we demonstrate a significantly tighter regret bound for VBOS, matching the strong regret bounds of Thompson sampling (Russo & Van Roy, 2014) and GP-UCB (Srinivas et al., 2010), regardless of the kernel.
@@ 4.1 Insights: Policy Initialization via Pre-Training & Context
The approximation error $D_{\sigma^{t}}(\pi^{t},\tilde{\pi}^{t})$ risks dominating the regret bound, since gradient-based optimization of the VBOS objective $\mathcal{V}$ yields a policy $\pi^{t}$ that lags behind the exact VBOS policy $\tilde{\pi}^{t}$. To mitigate this source of cumulative regret, $\pi^{t}$ must be initialized and maintained within a high-probability neighborhood of $\tilde{\pi}^{t}$, especially for large $t$, when $\tilde{\pi}^{t}$ tends to concentrate near the pre-trained policy. To support this, ToSFiT initializes $\pi^{t}$ via pre-training and prompt-conditioning, and adapts it cautiously toward the posterior PoM $\tilde{\pi}^{t}$ using small learning rates. These insights are illustrated in Figure 3 for a three-armed bandit. Note that Theorem 1 holds even if the pre-training prior $\pi^{0}$ is biased toward suboptimal regions. In that case, more compute effort in the form of gradient steps are required to ensure a vanishing $D_{\sigma^{t}}(\pi^{t},\tilde{\pi}^{t})$.
@@ 5 Experiments
We empirically evaluate ToSFiT across three diverse domains to assess its ability to (A) overcome the saturating performance of fixed generative foundation models without adaptation to posterior PoM, (B) leverage pre-training to accelerate optimization, and (C) scale effectively in batched settings. These experiments are designed to validate the theoretical insights from Section 4, particularly the importance of initializing from a strong prior and adapting cautiously toward the posterior PoM. We report the mean and standard error of the best-seen reward using $25$ random seeds, defined as the maximum reward observed up to step $t$, i.e., $\max_{t\leq n}r_{x^{t}}$. This metric captures the best solution found so far and is often used in black-box optimization benchmarks. Details are in Appendix D.

### v3

@@ Abstract
Bayesian optimization in large unstructured discrete spaces is often hindered by the computational cost of maximizing acquisition functions due to the absence of gradients. We propose a scalable alternative based on Thompson sampling that eliminates the need for acquisition function maximization by directly parameterizing the probability that a candidate yields the maximum reward. Our approach, Thompson Sampling via Fine-Tuning (ToSFiT) leverages the prior knowledge embedded in prompt-conditioned large language models, and incrementally adapts them toward the posterior. Theoretically, we derive a novel regret bound for a variational formulation of Thompson Sampling that matches the strong guarantees of its standard counterpart. Our analysis reveals the critical role of careful adaptation to the posterior probability of maximality—a principle that underpins our ToSFiT algorithm. Empirically, we validate our method on three diverse tasks: FAQ response refinement, thermally stable protein search, and quantum circuit design. Within a collection of methods covering Bayesian optimization, reinforcement learning, and evolutionary search, ToSFiT exhibits both state-of-the-art sample efficiency and computational efficiency.
@@ 1 Introduction
Bayesian optimization (Kushner, 1964; Garnett, 2023) is one such algorithmic framework that leverages Bayesian inference for large-scale experimental design and automated discovery, particularly in settings where experimental evaluations are costly or time-consuming. As a strategy for optimizing expensive black-box reward functions, it maintains a posterior distribution over the unknown rewards—typically modeled as a Gaussian process over a domain XX—and uses this model to guide the search for promising configurations. New candidates are selected by maximizing an acquisition function that balances two objectives: exploring uncertain regions to gather new information, and exploiting areas that are already known to perform well (Kushner, 1964; Auer, 2002).
Figure 1: ToSFiT treats candidate generation as Thompson sampling and fine-tunes the LLM toward the posterior PoM.
Among existing acquisition strategies, Thompson sampling (Thompson, 1933; Russo et al., 2018) stands out due to its state-of-the-art convergence guarantees (Russo and Van Roy, 2014; 2016; Chowdhury and Gopalan, 2017) and its strong empirical performance (Chapelle and Li, 2011), despite being one of the earliest strategies for Bayesian optimization. Thompson sampling draws a function from the reward posterior, which effectively serves as an acquisition function, and selects the point that maximizes it. The resulting evaluation points are distributed according to the probability of maximality (PoM) of rewards (Menet et al., 2025).
@@
In this work, we scale Bayesian optimization to large unstructured discrete spaces by instantiating Thompson Sampling as fine-tuning of large language models (ToSFiT, see Figure 1). Our approach builds on Variational Bayesian Optimistic Sampling (VBOS), established by O’Donoghue and Lattimore (2021), but introduces a key distinction: instead of optimizing toward the posterior probability of maximality starting from a uniform policy, we initialize it with a prompt-conditioned pre-trained LLM, and gradually adapt it toward the posterior. By treating generated proposals as Thompson samples, ToSFiT avoids intractable acquisition function maximization. Our contributions are:
@@
    We validate our method on FAQ response refinement, protein optimization, and quantum circuit design. Within a collection of methods covering Bayesian optimization, reinforcement learning, as well as evolutionary search, ToSFiT exhibits state-of-the-art performance.
@@ 2 Preliminaries
Consider an unknown reward vector $R$ over a discrete domain $X$. In Bayesian optimization, we want to find the $x\in X$ that maximizes $R_{x}$. Typically, we model our prior belief as Gaussian, i.e., $R\sim\mathcal{N}(\mu,K)$ with mean $\mu$ and covariance $K$. Thompson sampling suggests conducting Bayesian optimization by repeatedly conditioning the Gaussian reward vector $R$ on a noisy observation $Y_{x}:=R_{x}+\mathcal{N}(0,\sigma_{n}^{2})$ at a point $x\in X$ sampled from the posterior probability of maximality
@@
In practice, sampling from the probability of maximality is typically implemented by drawing a realization from the reward posterior and then selecting the point that maximizes this sample—effectively treating the realization as an acquisition function (Russo et al. 2018, see Figure 2). However, in large unstructured discrete domains, such maximization becomes intractable. To address this, we propose to directly parameterize the probability of maximality using a generative LLM. This approach, illustrated in Figure 1, avoids explicit maximization and, following the framework of O’Donoghue and Lattimore (2021), ensures consistency with the underlying reward model and thus vanishing regret (see Theorem 1).
@@ 2.2 Regret Bounds for Bayesian Optimization
Beyond the Bayesian setting, regret guarantees can also be established for the model-agnostic setting that assumes the reward $r$ to lie in a reproducing kernel Hilbert space $\mathbb{H}$ induced by a kernel $K:X\times X\to\mathbb{R}_{+}$. [^1] The expressivity of a reproducing kernel Hilbert space can vary widely with the choice of its kernel. For instance, the linear kernel induces the reproducing kernel Hilbert space of (all) linear functions, whereas the radial basis function kernel corresponds to a smooth subset of the square-integrable functions $L_{2}$ that can uniformly approximate any continuous function on a compact set. In this case, a Gaussian reward model is still used for inference, but its amplitude (i.e., prior uncertainty) must be chosen large enough to accommodate the complexity of $r$, as measured by the norm $\|r\|_{K}$. Notable examples of such model-agnostic algorithms include GP-UCB (Srinivas et al., 2010) and GP-TS (Chowdhury and Gopalan, 2017), a variant of Thompson sampling.
@@ 2.3 Variational Bayesian Optimistic Sampling
As shown in O’Donoghue and Lattimore (2021) and Tarbouriech et al. (2024), Gaussian PoM is effectively approximated by the VBOS policy $\tilde{\pi}$, the distribution that maximizes the functional
@@
Here, the reward surrogate $\hat{r}_{x}:=\mu_{x}+\sqrt{-2\ln({\pi_{x}})}\cdot\sigma_{x}$ represents a high-probability upper bound on $R_{x}$ (O’Donoghue and Lattimore, 2021). VBOS reveals an intimate connection (Tarbouriech et al., 2024; Menet et al., 2025) between PoM and upper confidence bounds augmented with entropy regularization (Ziebart, 2010; Haarnoja et al., 2018). Moreover, $\max_{\pi}\mathcal{V}(\pi)\geq\mathcal{V}(\mathrm{PoM})\geq\mathbb{E}[R^{*}]$ (see Corollary 2 in Appendix E), which is the key property that allows bounding the expected cumulative regret by $\tilde{\mathcal{O}}(\sqrt{T})$ for any optimistic policy $\pi$ with $\mathcal{V}(\pi)\geq\mathbb{E}[R^{*}]$. In addition to its variational form, Menet et al. (2025) pointed out a near-closed form with almost-linear runtime in $|X|$, given by
@@ 3 Thompson Sampling via Fine-Tuning
Algorithm 1 ToSFiT with Gaussian Process Reward Model
@@ 3.1 Scaling Gaussian Processes
The Moore–Aronszajn theorem states that any kernel $K$ can be expressed as an inner product in a reproducing kernel Hilbert space $\mathbb{H}$, i.e., $K(x,z)=\langle\phi(x),\phi(z)\rangle_{\mathbb{H}}$ for a feature map $\phi:X\to\mathbb{H}$. Thus we assume, without loss of generality, a linear kernel in $\mathbb{H}$, enabling scalable Gaussian process inference. Flexibility is retained through the choice of feature map $\phi$, which can be obtained from fixed embedding models (Ranković and Schwaller, 2023; Wei et al., 2022), learned adaptively during optimization (Ranković and Schwaller, 2025), or designed for the task at hand (Greenhill et al., 2020).
Furthermore, as detailed in Appendix D, we leverage a formulation of linear Gaussian processes that enables conditioning on observations, computing the reward posterior, and performing marginal likelihood maximization—all in closed form. As a result, both the computational and memory complexity scale in $\Theta(\text{dim}(\mathbb{H})^{2})$, independently of the number of observations, and has negligible overhead (see Table 2 in Section A.1 of the Appendix).
3.2 Gradients of Variational Bayesian Optimistic Sampling
We derive explicit gradients of the VBOS objective, offering new insights into its optimization landscape. All proofs are in Appendix E. Note that the score trick (Williams, 1992) cannot be applied blindly, since the reward surrogate $\hat{r}_{x}:=\mu_{x}+\sqrt{-2\ln({\pi_{x}^{\theta}})}\cdot\sigma_{x}$ depends on the policy $\pi^{\theta}$.
@@ 3.2 Gradients of Variational Bayesian Optimistic Sampling
$\tfrac{d}{d\theta}\mathcal{V}(\pi^{\theta})=\mathbb{E}_{x\sim\pi^{\theta}}\big[\big(\mu_{x}\underbrace{-\xi-v^{\text{-}1}(\pi^{\theta}_{x})\cdot\sigma_{x}}_{-\mu_{x}^{\theta}\text{ for }\xi=\kappa}\big)\cdot\tfrac{d}{d\theta}\ln\pi^{\theta}_{x}\big].$
@@ 3.3 Stabilizing Gradients of VBOS
Here, $\hat{\hat{r}}^{\theta}_{x_{i}}-\xi_{i}$ is referred to as the advantage function. In practice, this estimator can suffer from high variance, depending on the choice of baselines $\xi_{i}\in\mathbb{R}$. To address this, we adopt the Reinforce Leave-One-Out (RLOO) baseline (Kool et al., 2019), where each $\xi_{i}$ is set to the average of the other surrogates in the batch: $\xi_{i}=\tfrac{1}{B-1}\sum_{j\not=i}\hat{\hat{r}}^{\theta}_{x_{j}}$. This technique has been demonstrated to outperform more complex alternatives like Proximal Policy Optimization (Schulman et al., 2017) in fine-tuning large language models (Ahmadian et al., 2024). To further stabilize learning, we normalize the advantage function by its empirical standard deviation, derived from the empirical second moments: $\widehat{\text{advantage std}}=\sqrt{\tfrac{1}{B}\sum_{h}(\hat{\hat{r}}^{\theta}_{x_{h}}-\xi_{h})^{2}}$, which amounts to a variance-adaptive learning rate. As shown in Proposition 3 of Appendix C, standardized RLOO is mathematically equivalent to the advantage function used in Group Relative Policy Optimization (Shao et al., 2024).
@@ 4 Theory
Even the exact VBOS policy $\tilde{\pi}$ only conducts approximate Thompson sampling. Nevertheless, O’Donoghue and Lattimore (2021) prove an upper bound on the cumulative regret incurred when sampling a bandit according to exact VBOS, i.e., for $x^{t}\!\sim\!\tilde{\pi}^{t}$. The bound reads $\mathbb{E}[\sum_{t}R^{*}-R_{x^{t}}]\leq\sqrt{2|X|T\ln|X|(1\!+\!\ln T)}$. Note that the structure of the kernel is not taken into account, i.e., the worst-case bandit with independent arms is assumed. In Theorem 1, we demonstrate a significantly tighter regret bound for VBOS, matching the strong regret bounds of Thompson sampling (Russo and Van Roy, 2014) and GP-UCB (Srinivas et al., 2010), regardless of the kernel.
@@ 4.1 Insights: Policy Initialization via Pre-Training & Context
The approximation error $D_{\sigma^{t}}(\pi^{t},\tilde{\pi}^{t})$ risks dominating the regret bound, since gradient-based optimization of the VBOS objective $\mathcal{V}$ yields a policy $\pi^{t}$ that lags behind the exact VBOS policy $\tilde{\pi}^{t}$. To mitigate this source of cumulative regret, $\pi^{t}$ must be initialized and maintained within a high-probability neighborhood of $\tilde{\pi}^{t}$, especially for large $t$, when $\tilde{\pi}^{t}$ tends to concentrate near the pre-trained policy. To support this, ToSFiT initializes $\pi^{t}$ via pre-training and prompt-conditioning, and adapts it cautiously toward the posterior PoM $\tilde{\pi}^{t}$ using small learning rates. These insights are illustrated in Figure 3 for a three-armed bandit and experimentally verified in Section 5.6. Note that Theorem 1 holds even if the pre-training prior $\pi^{0}$ is biased toward suboptimal regions. In that case, more compute effort in the form of gradient steps are required to ensure a vanishing $D_{\sigma^{t}}(\pi^{t},\tilde{\pi}^{t})$.
@@ 5 Experiments
We empirically evaluate ToSFiT across three diverse domains for language models between 0.60.6B and 88B parameters. We report the mean and standard error of the best-seen reward using $25$ random seeds, defined as the maximum reward observed up to step $t$, i.e., $\max_{t\leq n}r_{x^{t}}$. This metric captures the best solution found so far and is often used in black-box optimization benchmarks. Details are in Appendix B. Within a collection of methods covering Bayesian optimization, reinforcement learning, and evolutionary search, ToSFiT exhibits state-of-the-art sample efficiency and computational efficiency. ToSFiT can naturally be applied in batched settings, and sample efficiency can be further improved by adjusting the computational effort per round of Bayesian optimization.
5.1 Settings
FAQ Refinement
is a natural language task that tests the algorithm’s ability to optimize text based on semantic alignment. We ask a Qwen3-1.7B/Qwen3-8B model (Yang et al., 2025) to write a frequently-asked-questions (FAQ) response. The reward is modeled as the alignment to an unknown ground-truth, judged by the Qwen3-Embedding-0.6B model (Zhang et al., 2025). As a deep kernel, we adopt a linear kernel over the first 256 entries of the embeddings of Qwen3-Embedding-0.6B. The search space consists of all token sequences, which is exponentially large in the response length.
Protein Search
explores the challenge of designing thermally stable proteins—a task with significant implications for drug development and industrial biotechnology. The goal is to identify amino acid sequences that exhibit high thermal stability, a property that enhances protein robustness and shelf life. Note that with 20 standard amino acids in the human body and sequence lengths of $100$ and above, the search space exceeds the number of atoms in the observable universe. We sample amino acid sequences from ProtGPT2 (Ferruz et al., 2022) (0.738B parameters) and score them according to their negative thermal instability index (Guruprasad et al., 1990). Two baselines, FIBO and Evolutionary Search (LLM), require instruction-tuned models. There, we use a Qwen3-0.6B/Qwen3-8B model. To predict the thermal stability and assess epistemic uncertainty, we adopt a GP with linear kernel over the mean token embeddings from ProtGPT2 projected to the unit sphere.
Quantum Circuit Design
is the task of designing quantum circuits that prepare low-energy quantum states in unknown environments. The challenge lies in navigating a vast, discrete space of valid quantum programs, where entanglement and gate structure critically influence performance. To generate Qiskit circuits (Javadi-Abhari et al., 2024), we use a Qwen2.5-Coder-1.5B/Qwen2.5-Coder-7B model (Hui et al., 2024) prompted to fill-in-the-middle (Bavarian et al., 2022) after initializing six disentangled qubits in the zero state. The baselines FIBO and Evolutionary Search (LLM) require instruction-tuned models, for which we use Qwen2.5-Coder-1.5B-Instruct/Qwen2.5-Coder-7B-Instruct. As reward, we consider the negative energy of the prepared state under an unknown Hamiltonian with strong interaction terms, requiring entanglement for optimal performance. The feature map for the linear Gaussian process over rewards consists of a code validity bit as well as all two-qubit Pauli observables, which can be efficiently simulated using quantum computers.
(a) FAQ Response Refinement
using Qwen3-1.7B
(b) Protein Search
using ProtGPT2-0.7B or Qwen3-0.6B
(c) Quantum Circuit Design
using Qwen2.5-Coder-1.5B(-Instruct)
(d) FAQ Response Refinement
using Qwen3-8B
(e) Protein Search
using ProtGPT2-0.7B or Qwen3-8B
(f) Quantum Circuit Design
using Qwen2.5-Coder-7B(-Instruct)
Figure 4: Across three tasks and within a collection of methods covering Bayesian optimization, reinforcement learning, as well as evolutionary search, ToSFiT exhibits state-of-the-art performance. In (b) and (e) Evolutionary Search (LLM) and FIBO use Qwen for instruction following.